find an equation of a line passing through (-2,-1) and parallel to y=-3x-1. show work
Answer by mananth(16946)

1 y = -3 x -1
Divide by 1
y = -3 x -1
Compare this equation with y=mx+b
slope m = -3
The slope of a line parallel to the above line will be the same
The slope of the required line will be -3
m= -3 ,point ( -2 , -1 )
Find b by plugging the values of m & the point in
y=mx+b
-1 = 6 + b
b= -7
m= -3
Plug value of the slope and b in y = mx +b
The required equation is y = -3 x -7

m.ananth@hotmail.ca